1. Suppose country A is a small import country (price taker) in the world market of lamp trade. Its domestic demand on and supply for lamp are: D=18-P, S=-12+2P. The world market price of lamp is 8. 

(1)  Please calculate the equilibrium price and quantity of lamp in country A before trade.

(2)  Please calculate A’s quantity of imported lamp under free trade.

Please draw graph and calculate the net welfare effect of free trade on A’s consumers, producers and the whole country.


1
Expert's answer
2021-09-12T19:31:34-0400

a)

"QP=18-P\\QS=-12+2P\\Qd=Qs\\18-P=-12+2P\\18+12=3P\\P=10\\Q=18-10\\Q=8units"


b)

World price 8

Required domestic

"Q's=-12+2(8)=4units\\Qd=18-8=10units\\Qs>Qd, surplus"

There will be export of "10-4=10units"


c)

"Consumer\space surplus\\0=18-P\\P=18\\\frac{1}{2}×(18-8)×10\\=\$50"

"Producer\space surplus\\0=-12-2P\\2P=12\\P=6\\\frac{1}{2}×(8-6)×10\\=\$10"